Budgeted Income Statement and Supporting Budgets The budget director of Birds of a Feather Inc., with the assistance o...

  1. Budgeted Income Statement and Supporting Budgets

    The budget director of Birds of a Feather Inc., with the assistance of the controller, treasurer, production manager, and sales manager, has gathered the following data for use in developing the budgeted income statement for January:

    1. Estimated sales for January:
        Birdhouse 6,000 units at $55 per unit
        Bird feeder 4,500 units at $75 per unit
    2. Estimated inventories at January 1:
      Direct materials:
        Wood 220 ft.
        Plastic 250 lb.
      Finished products:
        Birdhouse 300 units at $23 per unit
        Bird feeder 240 units at $34 per unit
    3. Desired inventories at January 31:
      Direct materials:
        Wood 180 ft.
        Plastic 210 lb.
      Finished products:
        Birdhouse 340 units at $23 per unit
        Bird feeder 200 units at $34 per unit
    4. Direct materials used in production:
      In manufacture of BirdHouse:
        Wood 0.80 ft. per unit of product
        Plastic 0.50 lb. per unit of product
      In manufacture of Bird Feeder:
        Wood 1.20 ft. per unit of product
        Plastic 0.75 lb. per unit of product
    5. Anticipated cost of purchases and beginning and ending inventory of direct materials:
        Wood $8.00 per ft.
        Plastic $1.20 per lb.
    6. Direct labor requirements:
      Birdhouse:
        Fabrication Department 0.20 hr. at $15 per hr.
        Assembly Department 0.30 hr. at $12 per hr.
      Bird Feeder:
        Fabrication Department 0.40 hr. at $15 per hr.
        Assembly Department 0.35 hr. at $12 per hr.
    7. Estimated factory overhead costs for January:
      Indirect factory wages $80,000
      Depreciation of plant and equipment 25,000
      Power and light 8,000
      Insurance and property tax 2,000
    8. Estimated operating expenses for January:
      Sales salaries expense $90,000
      Advertising expense 20,000
      Office salaries expense 18,000
      Depreciation expense—office equipment 800
      Telephone expense—selling 500
      Telephone expense—administrative 200
      Travel expense—selling 5,000
      Office supplies expense 250
      Miscellaneous administrative expense 450
    9. Estimated other income and expense for January:
      Interest revenue $300
      Interest expense 224
    10. Estimated tax rate: 30%

    Required:

    1. Prepare a sales budget for January.

    Birds of a Feather Inc.
    Sales Budget
    For the Month Ending January 31
    Unit Sales
    Volume
    Unit Selling
    Price
    Total Sales
    Birdhouse
    Bird feeder
    Total revenue from sales $

    2. Prepare a production budget for January.

    Birds of a Feather Inc.
    Production Budget
    For the Month Ending January 31
    Units
    Birdhouse Bird Feeder
    Expected units to be sold
    Plus desired inventory, January 31
      Total
    Less estimated inventory, January 1
    Total units to be produced

    3. Prepare a direct materials purchases budget for January.

    Birds of a Feather Inc.
    Direct Materials Purchases Budget
    For the Month Ending January 31
    Wood Plastic Total
    Required units for production:
      Birdhouse
      Bird feeder
    Plus desired units of inventory, January 31
    Total
    Less estimated units of inventory, January 1
    Total units to be purchased
    Unit price $ $
    Total direct materials to be purchased $ $ $

    4. Prepare a direct labor cost budget for January.

    Birds of a Feather Inc.
    Direct Labor Cost Budget
    For the Month Ending January 31
    Fabrication
    Department
    Assembly Department Total
    Hours required for production:
    Birdhouse
    Bird feeder
    Total
    Hourly rate $ $
    Total direct labor cost $ $ $

    5. Prepare a factory overhead cost budget for January.

    Birds of a Feather Inc.
    Factory Overhead Cost Budget
    For the Month Ending January 31
    Indirect factory wages
    Depreciation of plant and equipment
    Power and light
    Insurance and property tax
    Total $

    6. Prepare a cost of goods sold budget for January. Work in process at the beginning of January is estimated to be $29,000, and work in process at the end of January is estimated to be $35,400.

    Birds of a Feather Inc.
    Cost of Goods Sold Budget
    For the Month Ending January 31
    Direct materials:
       
       
      Cost of direct materials available for use
       
      Cost of direct materials placed in production
    Total manufacturing costs
    Total work in process during the period
    Cost of goods manufactured
    Cost of finished goods available for sale
    Cost of goods sold $

    7. Prepare a selling and administrative expenses budget for January.

    Birds of a Feather Inc.
    Selling and Administrative Expenses Budget
    For the Month Ending January 31
    Selling expenses:
    Sales salaries expense
    Advertising expense
    Telephone expense—selling
    Travel expense—selling
    Total selling expenses
    Administrative expenses:
    Office salaries expense
    Depreciation expense—office equipment
    Telephone expense—administrative
    Office supplies expense
    Miscellaneous administrative expense
    Total administrative expenses
    Total operating expenses $

    8. Prepare a budgeted income statement for January.

    Birds of a Feather Inc.
    Budgeted Income Statement
    For the Month Ending January 31
    Selling and administrative expenses:
    Total selling and administrative expenses
    Other revenue:
    Other expenses:
    $
Answer #1
Sales Budget
January
Birdhouse Birdfeeder Total
Unit Sales Volume 6000 4500 10500
Unit Selling Price $                55.00 $                75.00
Total Revenue from sales $     330,000.00 $     337,500.00 $ 667,500.00
Production Budget
January
Birdhouse Birdfeeder Total
Expected Unit Sales 6000 4500 10500
Add : Desired Ending Finished Good Units 340 200
Total Required Units 6340 4700 11040
Less : Beginning Finished Good units 300 240
Required Production units 6040 4460 10500
Direct Materials Budget
January
Wood Plastic Total
Required Units for production
Birdhouse 4832 =6040*0.8 3020 =6040*0.5 7852
Birdfeeder 5352 =4460*1.2 3345 =4460*0.75 8697
Add : Desired Ending inventory 180 210
Total Required Units 10364 6575 16939
Less : Beginning inventory 220 250
Total Units to be Purchased 10144 6325 16469
Unit Price $                  8.00 $               1.20
Total costs for direct materials $       81,152.00 $      7,590.00 $ 88,742.00
Direct Labor Budget
January
Fabrication
Department
Assembly
Department
Total
Hours Required for Production
Birdhouse 1208 =6040*0.2 1812 =6040*0.3 3020
Birdfeeder 1784 =4460*0.4 1561 =4460*0.35 3345
Total hours 2992 3373 6365
Hourly Rate $                15.00 $            12.00
Total cost of direct labor $       44,880.00 $    40,476.00 $ 85,356.00